# CAMPBELL & REMIRO-AZÓCAR

### v1

@@ Acknowledgments
Figure 3: Density plots showing the overlap of covariates $X_{1}$,$X_{2}$,$X_{3}$, and $X_{4}$ for Scenarios KS1 and KS2 in the simulation study.
Figure 4: Density plots showing the overlap of covariates $X_{1}$,$X_{2}$,$X_{3}$, and $X_{4}$ for Scenarios KS3 and KS4 in the simulation study.
@@
library(survminer)
library(flextable)
library(officer)
# set seed to give reproducible example
@@
# Read in ADaM data and rename variables of interest
@@
adtte <- read.csv(system.file("extdata", "adtte.csv", package = "MAIC", mustWork = TRUE))
adsl <- adsl %># Data containing the matching variables
  mutate(SEX=ifelse(SEX=="Male", 1, 0)) # Coded 1 for males and 0 for females
adrs <- adrs %># Response data
@@
adtte <- adtte %># Time to event data (overall survival)
  filter(PARAMCD=="OS") %>%
  mutate(Event=1-CNSR) %>#Set up coding as Event = 1, Censor = 0
  transmute(USUBJID, ARM, Time=AVAL, Event)
# Combine all intervention data
@@
  full_join(adrs, by=c("USUBJID", "ARM")) %>%
  full_join(adtte, by=c("USUBJID", "ARM"))
head(intervention_input)
# List out matching covariates
match_cov <- c("AGE",
               "SEX",
               "SMOKE",
               "ECOG0")
@@
# Renames target population cols to be consistent with match_cov
match_cov
names(target_pop)
target_pop_standard <- target_pop %>%
  #EDIT
  dplyr::rename(N=N,
         Treatment=ARM,
         AGE=age.mean,
         SEX=prop.male,
         SMOKE=prop.smoke,
         ECOG0=prop.ecog0
  transmute(N, Treatment, AGE, SEX, SMOKE, ECOG0)
target_pop_standard
@@
comparator_n <- target_pop\$N # total number of patients in the comparator data
comparator_prop_events <- 0.4 # proportion of responders
# Calculate number with event
# Use round() to ensure we end up with a whole number of people
# number without an event = Total N - number with event to ensure we keep the same number of patients
@@
comparator_binary <- data.frame("response"= c(rep(1, n_with_event), rep(0, comparator_n - n_with_event)))
n0 <- dim(comparator_binary)[1]
n1 <- length(intervention_input\$response)
Y_all <- unlist(c(intervention_input\$response,comparator_binary))
@@
X_all <- rbind(cbind(intervention_input%>%select(AGE, SEX, SMOKE, ECOG0)),
               cbind(AGE=rep(NA, n0),
                     SEX=rep(NA, n0),
                     SMOKE=rep(NA, n0),
                     ECOG0=rep(NA, n0)))
@@
  ATC_naive =  g_function(mu1_naive) - g_function(mu0_naive)
  ATC_naive
  SE_g_mu1 <- sqrt(1/(length((Y_all[S_all==1]))*mu1_naive*(1-mu1_naive)))
  SE_g_mu0 <- sqrt(1/(length((Y_all[S_all==0]))*mu0_naive*(1-mu0_naive)))
  (sqrt((mu0_naive*(1-mu0_naive)/length((Y_all[S_all==0])))))
  naive_function <- function(data, indices) {
    temp <- data[indices,]
    return( g_function(mean(temp[temp[,"S_all"]==1, "Y_all"])))}
  boot_samples <- boot(data=data.frame(Y_all=Y_all, S_all=S_all),
                       statistic=naive_function, R=boot_n,
                       strata=S_all,
                       parallel = "multicore")
  SE_ATC <- sqrt(sd(boot_samples\$t,na.rm=TRUE)^2  + SE_g_mu0^2)
  SE_ATC
  ATC_naive_CI_boot<- c(ATC_naive-abs(qnorm(0.025))*SE_ATC, ATC_naive+abs(qnorm(0.025))*SE_ATC)
  round(ATC_naive_CI_boot,3)
  SE_ATC <- sqrt(SE_g_mu1^2  + SE_g_mu0^2)
  ATC_naive_CI<- c(ATC_naive-abs(qnorm(0.025))*SE_ATC, ATC_naive+abs(qnorm(0.025))*SE_ATC)
  round(ATC_naive_CI,3)
  round((c(ATC_naive, ATC_naive_CI)),3)
@@
  # 5.318 3.880 7.289
  # compare with published result:
  # 5.318 (3.888 to 7.275)  https://roche.github.io/MAIC/articles/MAIC.html
  # 3.3 Entropy balancing (matching-adjusted indirect comparison)
  # ...equivalent to minimizing the objective function:
  objfn <- function(a1, X){ sum(exp(X %*a1)) }
  gradfn <- function(a1, X){ colSums(sweep(X, 1, exp(X %*a1), "*")) }
@@
  library(maicChecks)
  maicLP(AC.IPD[,-1], BC.ALD) # Checks if AD is within the convex hull of IPD using lp-solve
  X.EM.0 <- sweep(cbind(as.matrix((AC.IPD[, cov_names]))), 2,
                  as.matrix((BC.ALD[,c(paste("mean.",cov_names, sep=""))])), ’-’)
  gamma <- optim(par = rep(0,dim(X.EM.0)[2]),
@@
  mu0_EB <- mu0_naive
  ATC_EB <- g_function(mu1_EB) - g_function(mu0_EB)
  ATC_EB
  hist(wt_EB, xlab="Entropy balancing weights", main="")
@@
    objfn <- function(a1, X){ sum(exp(X %*a1)) }
    gradfn <- function(a1, X){ colSums(sweep(X, 1, exp(X %*a1), "*")) }
@@
    X.EM.0 <- sweep(cbind(as.matrix((AC.IPD[, cov_names]))), 2,
                    as.matrix((BC.ALD[,c(paste("mean.",cov_names, sep=""))])), ’-’)
    gamma <- optim(par = rep(0,dim(X.EM.0)[2]),
@@
  sd(boot_samples\$t,na.rm=TRUE)
  SE_ATC <- sqrt(sd(boot_samples\$t,na.rm=TRUE)^2  + SE_g_mu0^2)
  SE_ATC
  ATC_EB_CI <- c(ATC_EB-abs(qnorm(0.025))*SE_ATC,ATC_EB+abs(qnorm(0.025))*SE_ATC)
  round((c(ATC_EB, ATC_EB_CI)),3)
@@
  # 3.787 2.497 5.742
  # compare with published result:
  # 3.787 (2.558 to 5.605)  https://roche.github.io/MAIC/articles/MAIC.html
@@
  out2<- add_integration(data.frame(Y_all=NA),
@@
  Y_all <-  c(Y_all[S_all==1], c(rep(1, n_with_event), rep(0, M - n_with_event)))
@@
  n <- n1 + n0
  # The G-computation estimator
  # The G-computation estimator for the ATC contrasts the average of potential counterfactual outcomes
  # under the active intervention with the average of observed outcomes for the external control.
@@
                       data = data.frame(y=c(Y_all[S_all==1]),
                                         x=(X_all[S_all==1,])),
                       family = binomial(link="logit"))Ψ
  Y1_hat <- (predict(outcome_model, newdata = data.frame(x=X_all[S_all==0,]),
                     type = "response"))
@@
  mu0_GCOMP1 <- mu0_naive
  ATC_GCOMP <-  g_function(mu1_GCOMP) - g_function(mu0_GCOMP1)
  ATC_GCOMP
@@
                         data = data.frame(y=c(temp[,"Y_all"][temp[,"S_all"]==1]),
@@
                         family = "binomial")Ψ
    Y1_hat <- (predict(outcome_model, newdata = data.frame(x=temp[temp[,"S_all"]==0,grep("X_all",colnames(temp))]),
                       type = "response"))
@@
  sd(boot_samples\$t,na.rm=TRUE)
  SE_ATC <- sqrt(sd(boot_samples\$t,na.rm=TRUE)^2  + SE_g_mu0^2)
  ATC_GCOMP_CI <- c(ATC_GCOMP-abs(qnorm(0.025))*SE_ATC,ATC_GCOMP+abs(qnorm(0.025))*SE_ATC)
  round((c(ATC_GCOMP, ATC_GCOMP_CI)),3)
@@
  # DR3
  # Our novel contribution is augmenting the entropy balancing-based MAIC approach with an outcome model
  data_for_outcome_model <- data.frame(y=c(Y_all[S_all==1]),
                                       X_all[S_all==1,])
@@
  outcome_model <- glm(y ~ .,
                       data = data_for_outcome_model,
                       family = "binomial")Ψ
  Y1_hat_all <- (predict(outcome_model, newdata = data.frame(X_all),
                         type = "response"))
@@
  mu0_DR3 <- mu0_naive
  ATC_DR3 <-  g_function(mu1_DR3) - g_function(mu0_DR3)
  ATC_DR3
@@
                         data = data.frame(y=c(temp[,"Y_all"][temp[,"S_all"]==1]),
@@
                         family = "binomial")Ψ
    Y1_hat_all <- (predict(outcome_model, newdata = data.frame(x=temp[,grep("X_all",colnames(temp))]),
                           type = "response"))
    AC.IPD <- data.frame(y=temp[,"Y_all"][temp[,"S_all"]==1], X= temp[temp[,"S_all"]==1,grep("X_all",colnames(temp))])
@@
    objfn <- function(a1, X){ sum(exp(X %*a1)) }
    gradfn <- function(a1, X){ colSums(sweep(X, 1, exp(X %*a1), "*")) }
@@
    X.EM.0 <- sweep(cbind(as.matrix((AC.IPD[, cov_names]))), 2,
                    as.matrix((BC.ALD[,c(paste("mean.",cov_names, sep=""))])), ’-’)
    gamma <- optim(par = rep(0,dim(X.EM.0)[2]),
@@
  sd(boot_samples\$t,na.rm=TRUE)
  SE_ATC <- sqrt(sd(boot_samples\$t,na.rm=TRUE)^2  + SE_g_mu0^2)
  SE_ATC
  ATC_DR3_CI <- c(ATC_DR3-abs(qnorm(0.025))*SE_ATC,ATC_DR3+abs(qnorm(0.025))*SE_ATC)
  ATC_DR3_CI
  round((c(ATC_DR3, ATC_DR3_CI)),3)

### v2

@@ Acknowledgments
AI disclosure
During the preparation of this work, the authors used AI tools (Claude) for editing purposes. The authors have reviewed and edited the final content.
@@
Targeting the average treatment effect in the treated
We briefly adapt the methodologies in Section 3.2 to Section 3.6 of the main text so that these target the ATT. We assume that there is full IPD availability and that $\hat{\mu}_{1}^{1}=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}Y_{i}$ is consistent for $\mu_{1}^{1}$.
For the modeling-based IOW approaches in Section 3.2, external control subjects are weighted by their conditional odds of SAT participation to transport the external control outcomes to the SAT (sub) population. SAT subjects are unweighted and external control subjects $i=n_{1}+1,\dots,n$ are weighted by $\hat{w}_{i}=\hat{e}_{i}/(1-\hat{e}_{i})$. Assuming correct specification of the propensity score model, the estimated weights would balance the covariate distribution of the external control with respect to that of the SAT, enabling consistent estimation of mean absolute outcome $\mu^{0}_{1}$ and the ATT. Propensity score predictions that are close to one lead to extreme weights and imprecise ATT estimation, particularly where the sample size of the external control is small.
[…]
$\widehat{\textrm{ATT}}=g\underbrace{\left(\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}Y_{i}\right)}_{\hat{\mu}_{1}^{1}}-g\underbrace{\left(\frac{1}{K}\sum_{i=n_{1}+1}^{n}\hat{u}_{i}\epsilon_{i}^{0}+\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\hat{Y_{i}}^{0}\right)}_{\hat{\mu}_{1}^{0}},$
where $K$ is a constant, $\hat{u}_{i}$ is a weight estimate and $\epsilon_{i}^{0}=Y_{i}-\hat{Y}_{i}^{0}$ is a residual term for subject $i=n_{1}+1,\dots,n$ in the external control.
A weighted G-computation estimator akin to that described in Section 3.6 would target the ATT by: (1) estimating weights for the odds of SAT participation; (2) fitting a weighted model $m({\mathbf{X}_{i}};\hat{\mathbf{\beta}}_{v})$ for the conditional outcome expectation to the external control participants; and (3) averaging the outcome predictions of the weighted regression over the SAT covariate distribution. The resulting estimator for the mean absolute outcome $\mu_{1}^{0}$ is $\hat{\mu}_{1}^{0}=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\hat{Y}_{i}^{0}=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}}_{v})\right)$, which is then substituted into Equation 36 for estimation of the ATT.
@@
Figure 3: Density plots showing the overlap of covariates $X_{1}$$X_{2}$$X_{3}$, and $X_{4}$ for Scenarios KS1 and KS2 in the simulation study.
Figure 4: Density plots showing the overlap of covariates $X_{1}$$X_{2}$$X_{3}$, and $X_{4}$ for Scenarios KS3 and KS4 in the simulation study.
@@
The following R code provides all calculations in the applied example:
@@
library(maicChecks)
@@
adsl <- adsl %>%
  mutate(SEX = ifelse(SEX == "Male", 1, 0))
adrs <- adrs %>%
@@
  full_join(adrs, by = c("USUBJID", "ARM"))
@@
comparator_n <- target_pop\$N
comparator_prop_events <- 0.4
@@
n0 <- comparator_n
n1 <- nrow(intervention_input)
Y_all <- c(intervention_input\$response, rep(1, n_with_event), rep(0, n0 - n_with_event))
@@
X_all <- rbind(
  intervention_input %>% select(AGE, SEX, SMOKE, ECOG0),
  data.frame(AGE = rep(NA, n0), SEX = rep(NA, n0),
             SMOKE = rep(NA, n0), ECOG0 = rep(NA, n0))
@@
ATC_naive <- g_function(mu1_naive) - g_function(mu0_naive)
SE_g_mu1 <- sqrt(1 / (n1 * mu1_naive * (1 - mu1_naive)))
SE_g_mu0 <- sqrt(1 / (n0 * mu0_naive * (1 - mu0_naive)))
SE_ATC_naive <- sqrt(SE_g_mu1^2 + SE_g_mu0^2)
ATC_naive_CI <- c(ATC_naive - qnorm(0.975) * SE_ATC_naive,
                  ATC_naive + qnorm(0.975) * SE_ATC_naive)
round(c(ATC_naive, ATC_naive_CI), 3)
@@
# Entropy balancing (MAIC)
objfn <- function(a1, X){ sum(exp(X %*% a1)) }
gradfn <- function(a1, X){ colSums(sweep(X, 1, exp(X %*% a1), "*")) }
@@
maicLP(AC.IPD[, -1], BC.ALD)
X.EM.0 <- sweep(as.matrix(AC.IPD[, cov_names]), 2,
                as.matrix(BC.ALD[, paste0("mean.", cov_names)]), ’-’)
gamma <- optim(par = rep(0, ncol(X.EM.0)),
@@
ATC_EB <- g_function(mu1_EB) - g_function(mu0_naive)
@@
  objfn <- function(a1, X){ sum(exp(X %*% a1)) }
  gradfn <- function(a1, X){ colSums(sweep(X, 1, exp(X %*% a1), "*")) }
@@
  X.EM.0 <- sweep(as.matrix(AC.IPD[, cov_names]), 2,
                  as.matrix(BC.ALD[, paste0("mean.", cov_names)]), ’-’)
  gamma <- optim(par = rep(0, ncol(X.EM.0)),
@@
SE_ATC_EB <- sqrt(sd(boot_samples\$t, na.rm = TRUE)^2 + SE_g_mu0^2)
ATC_EB_CI <- c(ATC_EB - qnorm(0.975) * SE_ATC_EB,
               ATC_EB + qnorm(0.975) * SE_ATC_EB)
round(c(ATC_EB, ATC_EB_CI), 3)
@@
out2 <- add_integration(
  data.frame(Y_all = NA),
@@
Y_all <- c(Y_all[S_all == 1], rep(1, n_with_event), rep(0, M - n_with_event))
@@
# G-computation estimator
@@
                     data = data.frame(y = Y_all[S_all == 1],
                                       x = X_all[S_all == 1, ]),
                     family = binomial(link = "logit"))
Y1_hat <- predict(outcome_model, newdata = data.frame(x = X_all[S_all == 0, ]),
                  type = "response")
@@
ATC_GCOMP <- g_function(mu1_GCOMP) - g_function(mu0_naive)
@@
                       data = data.frame(y = temp[, "Y_all"][temp[, "S_all"] == 1],
@@
                       family = "binomial")
  Y1_hat <- predict(outcome_model,
                    newdata = data.frame(x = temp[temp[, "S_all"] == 0, grep("X_all", colnames(temp))]),
                    type = "response")
@@
SE_ATC_GCOMP <- sqrt(sd(boot_samples\$t, na.rm = TRUE)^2 + SE_g_mu0^2)
ATC_GCOMP_CI <- c(ATC_GCOMP - qnorm(0.975) * SE_ATC_GCOMP,
                  ATC_GCOMP + qnorm(0.975) * SE_ATC_GCOMP)
round(c(ATC_GCOMP, ATC_GCOMP_CI), 3)
@@
# DR augmented MAIC estimator
data_for_outcome_model <- data.frame(y = Y_all[S_all == 1], X_all[S_all == 1, ])
@@
outcome_model <- glm(y ~ ., data = data_for_outcome_model, family = "binomial")
Y1_hat_all <- predict(outcome_model, newdata = data.frame(X_all), type = "response")
@@
ATC_DR3 <- g_function(mu1_DR3) - g_function(mu0_naive)
@@
                       data = data.frame(y = temp[, "Y_all"][temp[, "S_all"] == 1],
@@
                       family = "binomial")
  Y1_hat_all <- predict(outcome_model,
                        newdata = data.frame(x = temp[, grep("X_all", colnames(temp))]),
                        type = "response")
  AC.IPD <- data.frame(y = temp[, "Y_all"][temp[, "S_all"] == 1],
                       X = temp[temp[, "S_all"] == 1, grep("X_all", colnames(temp))])
@@
  objfn <- function(a1, X){ sum(exp(X %*% a1)) }
  gradfn <- function(a1, X){ colSums(sweep(X, 1, exp(X %*% a1), "*")) }
@@
  X.EM.0 <- sweep(as.matrix(AC.IPD[, cov_names]), 2,
                  as.matrix(BC.ALD[, paste0("mean.", cov_names)]), ’-’)
  gamma <- optim(par = rep(0, ncol(X.EM.0)),
@@
SE_ATC_DR3 <- sqrt(sd(boot_samples\$t, na.rm = TRUE)^2 + SE_g_mu0^2)
ATC_DR3_CI <- c(ATC_DR3 - qnorm(0.975) * SE_ATC_DR3,
                ATC_DR3 + qnorm(0.975) * SE_ATC_DR3)
round(c(ATC_DR3, ATC_DR3_CI), 3)
@@
# Normalized Inverse Odds Weighting (Hajek type)
ps_model <- glm(S ~ AGE + SEX + SMOKE + ECOG0 + AGE_SQ,
                data = data.frame(S = S_all, X_all),
                family = binomial(link = "logit"))
e_hat_intervention <- predict(ps_model, type = "response")[S_all == 1]
iow_raw <- (1 - e_hat_intervention) / e_hat_intervention
iow_normalized <- iow_raw / sum(iow_raw)
# Effective sample size
ESS_IOW <- (sum(iow_raw))^2 / sum(iow_raw^2)
# ATC estimation
mu1_IOW_Hajek <- sum(iow_normalized * intervention_input\$response)
ATC_IOW_Hajek <- g_function(mu1_IOW_Hajek) - g_function(mu0_naive)
# Bootstrap for SE
ATC_IOW_Hajek_function <- function(data, indices) {
  temp <- data[indices, ]
  ps_model_b <- glm(S ~ AGE + SEX + SMOKE + ECOG0 + AGE_SQ,
                    data = temp, family = binomial(link = "logit"))
  e_hat_b <- predict(ps_model_b, type = "response")
  e_hat_int_b <- e_hat_b[temp\$S == 1]
  iow_raw_b <- (1 - e_hat_int_b) / e_hat_int_b
  iow_norm_b <- iow_raw_b / sum(iow_raw_b)
  mu1_b <- sum(iow_norm_b * temp\$Y[temp\$S == 1])
  mu1_b <- max(min(mu1_b, 0.9999), 0.0001)
  return(g_function(mu1_b))
set.seed(123)
boot_samples_IOW <- boot(
  data = data.frame(Y = Y_all, S = S_all, X_all),
  statistic = ATC_IOW_Hajek_function,
  R = boot_n, strata = S_all, parallel = "multicore"
SE_ATC_IOW <- sqrt(sd(boot_samples_IOW\$t, na.rm = TRUE)^2 + SE_g_mu0^2)
ATC_IOW_Hajek_CI <- c(ATC_IOW_Hajek - qnorm(0.975) * SE_ATC_IOW,
                      ATC_IOW_Hajek + qnorm(0.975) * SE_ATC_IOW)
round(c(ATC_IOW_Hajek, ATC_IOW_Hajek_CI), 3)
round(exp(c(ATC_IOW_Hajek, ATC_IOW_Hajek_CI)), 3)
# Covariate balance table
fmt2 <- function(x) sprintf("%.2f", x)
# MAIC-weighted statistics
ESS_MAIC <- 1 / sum(wt_EB^2)
maic_mean_age   <- sum(wt_EB * intervention_input\$AGE)
maic_mean_sex   <- sum(wt_EB * intervention_input\$SEX)
maic_mean_ecog  <- sum(wt_EB * intervention_input\$ECOG0)
maic_mean_smoke <- sum(wt_EB * intervention_input\$SMOKE)
maic_sd_age     <- sqrt(sum(wt_EB * (intervention_input\$AGE - maic_mean_age)^2))
# IOW-weighted statistics
iow_mean_age   <- sum(iow_normalized * intervention_input\$AGE)
iow_mean_sex   <- sum(iow_normalized * intervention_input\$SEX)
iow_mean_ecog  <- sum(iow_normalized * intervention_input\$ECOG0)
iow_mean_smoke <- sum(iow_normalized * intervention_input\$SMOKE)
iow_sd_age     <- sqrt(sum(iow_normalized * (intervention_input\$AGE - iow_mean_age)^2))
balance_df <- data.frame(
  Covariate = c("Age in years (mean; SD)", "Sex (proportion male)",
                "ECOG (proportion status 1)", "Smoking (proportion smokers)"),
  Intervention = c(
    paste0(fmt2(mean(intervention_input\$AGE)), "; ", fmt2(sd(intervention_input\$AGE))),
    fmt2(mean(intervention_input\$SEX)),
    fmt2(mean(intervention_input\$ECOG0)),
    fmt2(mean(intervention_input\$SMOKE))),
  External = c(
    paste0(fmt2(target_pop\$age.mean), "; ", fmt2(target_pop\$age.sd)),
    fmt2(target_pop\$prop.male), fmt2(target_pop\$prop.ecog0), fmt2(target_pop\$prop.smoke)),
  IOW = c(
    paste0(fmt2(iow_mean_age), "; ", fmt2(iow_sd_age)),
    fmt2(iow_mean_sex), fmt2(iow_mean_ecog), fmt2(iow_mean_smoke)),
  MAIC = c(
    paste0(fmt2(maic_mean_age), "; ", fmt2(maic_sd_age)),
    fmt2(maic_mean_sex), fmt2(maic_mean_ecog), fmt2(maic_mean_smoke)),
  stringsAsFactors = FALSE
print(balance_df)
# Combined histogram of weights
weights_df <- data.frame(
  Weight = c(as.numeric(wt_EB), as.numeric(iow_normalized)),
  Method = factor(
    c(rep("MAIC (entropy balancing)", length(wt_EB)),
      rep("Normalized IOW", length(iow_normalized))),
    levels = c("Normalized IOW", "MAIC (entropy balancing)")
p_hist <- ggplot(weights_df, aes(x = Weight)) +
  geom_histogram(bins = 30, fill = "grey70", colour = "black", linewidth = 0.3) +
  facet_wrap(~ Method, scales = "free", ncol = 2) +
  labs(x = "Weight", y = "Frequency") +
  theme_minimal(base_size = 12) +
  theme(strip.text = element_text(face = "bold", size = 11),
        panel.grid.minor = element_blank())
print(p_hist)
# Forest plot comparing all estimators
forest_data <- data.frame(
  Estimator = factor(
    c("Naive", "IOW (Hajek)", "MAIC (EB)", "G-computation", "DR augmented MAIC"),
    levels = c("DR augmented MAIC", "G-computation", "MAIC (EB)", "IOW (Hajek)", "Naive")
  Estimate  = c(ATC_naive, ATC_IOW_Hajek, ATC_EB, ATC_GCOMP, ATC_DR3),
  CI_lower  = c(ATC_naive_CI[1], ATC_IOW_Hajek_CI[1], ATC_EB_CI[1], ATC_GCOMP_CI[1], ATC_DR3_CI[1]),
  CI_upper  = c(ATC_naive_CI[2], ATC_IOW_Hajek_CI[2], ATC_EB_CI[2], ATC_GCOMP_CI[2], ATC_DR3_CI[2])
forest_data\$label <- paste0(
  sprintf("%.3f", forest_data\$Estimate), " (",
  sprintf("%.3f", forest_data\$CI_lower), ", ",
  sprintf("%.3f", forest_data\$CI_upper), ")"
p_logOR <- ggplot(forest_data, aes(x = Estimate, y = Estimator)) +
  geom_vline(xintercept = 0, linetype = "dashed", colour = "grey50") +
  geom_errorbarh(aes(xmin = CI_lower, xmax = CI_upper),
                 height = 0.2, linewidth = 0.6) +
  geom_point(size = 4, shape = 20) +
  geom_text(aes(x = max(forest_data\$CI_upper) + 0.15, label = label),
            hjust = 0, size = 3.2) +
  labs(x = "ATC (marginal log-odds ratio)", y = NULL) +
  theme_minimal(base_size = 12) +
  theme(panel.grid.major.y = element_blank(),
        panel.grid.minor = element_blank(),
        axis.text.y = element_text(size = 11)) +
  coord_cartesian(xlim = c(min(forest_data\$CI_lower) - 0.1,
                           max(forest_data\$CI_upper) + 1.2))
print(p_logOR)